How many grams of Cu would have to react to produce 89.5 grams of Ag


1
Expert's answer
2022-03-19T02:44:52-0400

The reaction is:

"Cu + 2Ag(+)=>Cu(2+) +2Ag"


According to the equation, 0.5 mol (or 1 mol) of Cu would have to react to produce 1 mol of Ag (or 2 mol):

"\nu(Cu)=1/2\cdot\nu(Ag)"


"M(Cu)=63.5 g\/mol"

"M(Ag)=107.9 g\/mol"


89.5 g of Ag is:

"\nu(Ag)=m(Ag)/M(Ag)=89.5 g/(108g/mol)=0.83mol"



"\nu=m/M=>m=\nu\cdot M"

Then m(Cu):

"m(Cu)=\nu(Cu)\cdot M(Cu)=1/2\cdot\nu(Ag)\cdot M(Cu)=" "=0.5\cdot 0.83mol\cdot 63.5g/mol=26.3g"